OpenStudy (anonymous):

5x(2x^2-3x-1)=10x3-15x^2

OpenStudy (anonymous):

=5*2x^(2+1)-(5*3)x^(1+1)-(5*1)x=10x^3-15x^2-5x

OpenStudy (anonymous):

10x^3-15x^2-5x=10x^3-15x^2

OpenStudy (anonymous):

-5x=0 x=0

OpenStudy (anonymous):

i don't get it 😣

OpenStudy (anonymous):

ok the first step is to destribute 5x 5x(2x^2-3x-1) multiply it by every thing inside the parentheses

OpenStudy (anonymous):

nvm x does = 0

OpenStudy (anonymous):

you do 5x*2x^2-5x*3x-5x*1

OpenStudy (anonymous):

still confused 😔😣i suck at math

OpenStudy (anonymous):

then you simplify and get 10x^3-15x^2-5x=10x^3-15x^2 subtract 10x^3 from both sides get -15x^2-5x=-15x^2 add 15x^2 to both sides get -5x=0 divide by -5 x=0

OpenStudy (anonymous):

ohhh ok i feel dumb 😕

OpenStudy (anonymous):

its ok it is hard to learn don't feel dumb... people are here to help you